Phil's Carvings, Inc. wants to have a weighted average cost of capital of 9.5 percent. The firm has an aftertax cost of debt of 6.5 percent and a cost of equity of 12.75 percent. What debt-equity ratio is needed for the firm to achieve their targeted weighted average cost of capital?

a. .67

b. .84

c. .92

d. .76

e. 1.08
Answer #1

Let weight of Debt = x, then weight of Equity = 1-x.

So, Weighted average cost of capital,

WACC = (1-x)*12.75% + x*6.5% = 9.5%

By solving the above equation, we get x = 0.52

Hence, weight of Debt = 0.52 and weight of equity = 0.48

Debt equity ratio = weight of debt/weight of equity = 0.52/0.48 = 1.08

Option e is correct